OpenStudy (anonymous):

Solve for x: -3(x + 3) = -3(x + 1) - 5

OpenStudy (anonymous):

First remove brackets -3x -9 = -3x -3 -5 Then place like terms -9+3+5= -3x +3x then solve -1=0x so X = undefinable.